Find the rectangular equation for z=r^2 cos (2theta)

Name the surface.
1
Expert's answer
2020-03-10T14:34:25-0400

"z=r^2cos(2\theta)"

Solution:

"cos(2\theta)=cos^2\theta-sin^2\theta"

"z=r^2cos^2 \theta-r^2sin^2\theta"

Since, "x=rcos\theta, y=rsin\theta"

"z=x^2-y^2."

Answer:

Hyperbolic paraboloid: "z=x^2-y^2."